OpenStudy (anonymous):

Could someone explain Pythagorean triples using this question?: What whole number forms a Pythagorean triple with 40 and 41? a. 50 b. 9 c. 15 d. 39 Thanks! x

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

So any three numbers that "fit" into the Pythagorean theorem makes a Pythagorean triple?

OpenStudy (unklerhaukus):

exactly, a common example is the Pythagorean triple 3,4,5 because 3^2 + 4^2 = 5^2 9 + 16 = 25

OpenStudy (anonymous):

Got it. Thanks so much!

OpenStudy (anonymous):

So for this problem, what would the answer be? I just tried to work it out, and I must be doing something wrong because I'm not coming up with an answer that's a choice. I've also tried plugging in the choices….

OpenStudy (anonymous):

Just figured it out. :)

OpenStudy (unklerhaukus):

maybe you have the sides in the wrong order,

OpenStudy (anonymous):

Yup. That was my problem.

OpenStudy (unklerhaukus):

what final answer did you get?

OpenStudy (anonymous):

b. 9

OpenStudy (unklerhaukus):

YESssssss \[\color{red}\checkmark \]

OpenStudy (anonymous):

Haha thanks for your help!